# Fairness-Accuracy Trade-Offs: A Causal Perspective

### v1

@@ 1 Introduction
Automated decision-making systems based on machine learning and artificial intelligence are now commonly implemented in various critical sectors of society such as hiring, university admissions, law enforcement, credit assessments, and health care (Khandani et al., 2010; Mahoney and Mohen, 2007; Brennan et al., 2009). These technologies now significantly influence the lives of individuals and are frequently used in high-stakes settings (Topol, 2019; Berk et al., 2021; Taddeo and Floridi, 2018). As these systems replace or augment human decision-making processes, concerns about fairness and bias based on protected attributes such as race, gender, or religion have become a prominent consideration in the ML literature. The available data used to train automated systems may contain past and present societal biases as an imprint and therefore has the potential to perpetuate or even exacerbate discrimination against protected groups. This is highlighted by reports on biases in systems for sentencing Angwin et al. (2016), facial recognition Buolamwini and Gebru (2018), online ads Sweeney (2013); Datta et al. (2015), and system authentication Sanburn (2015), among many others. Despite the promise of AI to enhance human decision-making, the reality is that these technologies can also reflect or worsen societal inequalities. As alluded to before, the issue does not arise uniquely from the usage of automated systems; human-driven decision-making has long been analyzed in a similar fashion. Evidence of bias in human decision-making is abundant, including studies on the gender wage gap (Blau and Kahn, 1992, 2017) and racial disparities in legal outcomes (Sweeney and Haney, 1992; Pager, 2003). Therefore, without proper care about fairness and transparency of the new generation of AI systems, it is unclear what its impact will be on the historically discriminated groups, and issues of inequity more broadly.
@@
Within the growing literature on fair machine learning, a plethora of fairness definitions have been proposed. Commonly considered statistical criteria, among others, include demographic parity (independence (Darlington, 1971)), equalized odds (separation (Hardt et al., 2016)), and calibration (sufficiency (Chouldechova, 2017)). These definitions, however, have been shown as mutually incompatible (Barocas and Selbst, 2016; Kleinberg et al., 2016). Despite a number of proposals, there is still a lack of consensus on what the appropriate measures of fairness are, and how statistical notions of fairness could incorporate moral values of the society at large. For this reason, a number of works explored the causal approaches to fair machine learning (Kusner et al., 2017; Kilbertus et al., 2017; Nabi and Shpitser, 2018; Zhang and Bareinboim, 2018b, a; Wu et al., 2019; Chiappa, 2019; Plečko and Meinshausen, 2020), and an in-depth discussion can be found in (Plečko and Bareinboim, 2024). The main motivation for doing so is that the causal approach may allow the system designers to attribute the observed disparities between demographic groups to the causal mechanisms that underlie and generate them in the first place. In this way, by isolating disparities transmitted along different causal pathways, one obtains a more fine-grained analysis, and the capability to decide which causal pathways are deemed as unfair or discriminatory. More fundamentally, such considerations also form the basis of the legal frameworks for assessing discrimination in the United States and Europe. For instance, in the context of employment law, the disparate impact doctrine within the Title VII of the Civil Rights Act of 1964 (Act, 1964) disallows any form of discrimination that results in a too large of a disparity between groups of interest. A core aspect of this doctrine, however, is the notion of business necessity (BN) or job-relatedness. Considerations of business necessity may allow variables correlated with the protected attribute to act as a proxy, and the law does not necessarily prohibit their usage due to their relevance to the business itself (or more broadly the utility of the decision-maker). Often, the wording that is used is that to argue business necessity in front of a court of law, the plaintiff needs to demonstrate that “there is no practice that is less discriminatory and achieves the same utility” (Els, 1993). This concept illustrates the tension between fairness and utility, and demonstrates that we cannot be oblivious to considerations of utility from a legal standpoint.
@@
The question of fairness-utility trade-offs has been explored in the fairness literature (Corbett-Davies et al., 2017), with the essential argument that an unconstrained predictor always achieves a greater or equal utility than a constrained one. Despite this, the literature seems to be divided on this issue. For instance, some works argue that fairness and utility trade-offs are negligible in practice (Rodolfa et al., 2021), while others argue that such trade-offs need not even exist (Maity et al., 2020; Dutta et al., 2020). In this paper, we directly refute such arguments and demonstrate that, from a causal viewpoint, fairness and utility are always in a trade-off. The key subtlety in the works that find no trade-off (Rodolfa et al., 2021; Maity et al., 2020; Dutta et al., 2020) are the fairness metrics used to impose constraints. Naturally, the implications on the predictor’s utility will strongly depend on the exact type of the fairness constraint that is enforced [^1] Works that discuss how fairness-accuracy trade-offs do not exist or are insignificant in practice often focus on equality of odds (Hardt et al., 2016) ($\widehat{Y}{\perp\!\!\!\perp}X\mid Y$). Notably, this metric always allows for the perfect predictor $\widehat{Y}=Y$, and thus in settings with good predictive power, the cost of enforcing this constraint may indeed be negligible., and we demonstrate that imposing causal fairness constraints along different causal pathways almost always reduces predictive power.
@@
Interestingly, the tension between fairness and utility has been largely unexplored in the causal fairness literature (with some exceptions such as (Nilforoshan et al., 2022; Plecko and Bareinboim, 2024a; Plečko and Bareinboim, 2024)). Our main aim of this paper is to fill in this gap, and provide a systematic way of analyzing the fairness-accuracy trade-off from a causal lens. We illustrate our approach in a simple linear setting:
@@
Consider variables $X,W,Y$ behaving according to the following system of equations:
@@
where $\epsilon_{w}\sim N(0,\sigma_{w}^{2}),\epsilon_{y}\sim N(0,\sigma_{y}^{2})$. Variable $X$ is the protected attribute, and $Y$ is the outcome of interest. The causal diagram of Eqs. 1 -3 is shown in Fig. 3 (with the $Z$ set empty). Attribute $X$ can influence $Y$ along two different pathways: the direct path $X\to Y$, and the indirect path $X\to W\to Y$. Therefore, we consider fair predictors $\widehat{Y}^{S}$ of the form $\widehat{Y}^{S}=\hat{\alpha}_{S}X+\hat{\gamma}_{S}W,$ where the predictor $\widehat{Y}^{S}$ removes effects in the set $S$, with $S$ ranging in {,DE,IE,{DE, $,\{\text{DE, IE}\}\}$ , { DE, IE } } (DE, IE stand for direct and indirect effects). For instance, the optimal predictor $\widehat{Y}^{\emptyset}$ has $\hat{\alpha}_{\emptyset}=\alpha,\hat{\gamma}_{\emptyset}=\gamma$, and therefore its mean-squared error (MSE) equals $\mathbbm{E}[Y-\widehat{Y}^{\emptyset}]^{2}=\sigma_{y}^{2}$. The DE-fair predictor has $\hat{\alpha}_{\text{DE}}=0,\hat{\gamma}_{\text{DE}}=\gamma$, and the fully-fair predictor $\widehat{Y}^{\{\text{DE, IE}\}}$ has $\hat{\alpha}_{\{\text{DE, IE}\}}=0,\hat{\gamma}_{\{\text{DE, IE}\}}=0$. Thus, the corresponding MSE values can be written as (see Appendix A for details):
@@
$\mathbbm{E}[Y-\widehat{Y}^{\text{DE}}]^{2}=\sigma_{y}^{2}+\frac{\alpha^{2}}{2},\quad\mathbbm{E}[Y-\widehat{Y}^{\{\text{DE, IE}\}}]^{2}$ $=\sigma_{y}^{2}+\frac{\alpha^{2}+\gamma^{2}\beta^{2}+\alpha\gamma\beta}{2}+\gamma^{2}\sigma_{w}^{2}.$ (4)
@@
Our goal is to decompose the total excess loss originating from imposing the fairness constraints into the path-specific contributions:
@@
$\underbrace{\mathbbm{E}[Y-\widehat{Y}^{\{\text{DE, IE}\}}]^{2}}_{\text{fully-fair predictor's loss}}-\underbrace{\mathbbm{E}[Y-\widehat{Y}^{\emptyset}]^{2}}_{\text{unconstrained loss}}$ $=\underbrace{\frac{\alpha^{2}}{2}}_{\text{Term I}}+\underbrace{\frac{\gamma^{2}\beta^{2}+\alpha\gamma\beta}{2}+\gamma^{2}\sigma_{w}^{2}}_{\text{Term II}}$ (5)
@@
$=\underbrace{\mathbbm{E}[Y-\widehat{Y}^{\text{DE}}]^{2}-\mathbbm{E}[Y-\widehat{Y}^{\emptyset}]^{2}}_{\text{Term I = excess DE loss}}$ (6)
$+\underbrace{\mathbbm{E}[Y-\widehat{Y}^{\{\text{DE, IE}\}}]^{2}-\mathbbm{E}[Y-\widehat{Y}^{\text{DE}}]^{2}}_{\text{Term II = excess IE loss}}.$ (7)
At the same time, for each pathway, we can also observe the decrease in group disparity associated with removing the effect along the path $S^{\prime}$, by computing
$\underbrace{(\mathbbm{E}[\widehat{Y}^{\{\text{DE, IE}\}}\mid x_{1}]-\mathbbm{E}[\widehat{Y}^{\{\text{DE, IE}\}}\mid x_{0}])}_{\text{disparity after removing DE, IE}}-\underbrace{(\mathbbm{E}[Y^{\emptyset}\mid x_{0}]-\mathbbm{E}[Y^{\emptyset}\mid x_{0}])}_{\text{disparity before removing DE, IE}}=-\alpha-\beta\gamma,$ (8)
which again decomposes into the contributions along direct and indirect effects:
$(\mathbbm{E}[\widehat{Y}^{\text{DE}}\mid x_{0}]-\mathbbm{E}[\widehat{Y}^{\text{DE}}\mid x_{0}])-(\mathbbm{E}[\widehat{Y}^{\emptyset}\mid x_{0}]-\mathbbm{E}[\widehat{Y}^{\emptyset}\mid x_{0}])$ $=-\alpha$ (9)
$(\mathbbm{E}[\widehat{Y}^{\{\text{DE, IE}\}}\mid x_{1}]-\mathbbm{E}[\widehat{Y}^{\{\text{DE, IE}\}}\mid x_{0}])-(\mathbbm{E}[\widehat{Y}^{\text{DE}}\mid x_{0}]-\mathbbm{E}[\widehat{Y}^{\text{DE}}\mid x_{0}])$ $=-\beta\gamma.$ (10)
Prototypical fairness/utility trade-offs for predictors $\widehat{Y}^{S}$ are visualized in Fig. 1, for three randomly drawn triples $(\alpha,\beta,\gamma)$. Predictor $\widehat{Y}^{\emptyset}$ is optimal and thus always has 00 excess loss (and lies on the vertical axis). Fully-fair predictor $\widehat{Y}^{\{\text{DE, IE}\}}$ removes both direct and indirect effects, and thus always has 00 TV measure (and lies on the vertical axis). The slopes in the plot between $\widehat{Y}^{\emptyset}$, $\widehat{Y}^{\text{DE}}$, and $\widehat{Y}^{\text{DE, IE}}$ (indicated by arrows), geometrically capture the tension between excess loss and reducing discrimination upon imposing a constraint. These slopes are computed as TV differenceExcess $\frac{\text{TV difference}}{\text{Excess Loss}}$, a quantity that we call the causal fairness-utility ratio (CFUR). Based on Eqs. 6-7 and 9-10 we can compute that
@@
$\text{CFUR}(\text{DE})=-\frac{1}{\alpha},\quad\text{CFUR}(\text{IE})=-\frac{2\beta\gamma}{\gamma^{2}\beta^{2}+\alpha\gamma\beta+2\gamma^{2}\sigma_{w}^{2}},$ (11)
summarizing the fairness-utility trade-off for each pathway.
Figure 1: TV vs. Excess Loss.
@@
The above example illustrates how in the simple linear case we can attribute the increased loss from imposing fairness constraints to the specific causal pathway in question. It also shows we can compute the associated change in the disparity between groups, measured by the so-called TV measure $\mathbbm{E}[\widehat{Y}\mid x_{1}]-\mathbbm{E}[\widehat{Y}\mid x_{0}]$ (also known as the parity gap), a result known under the rubric of causal decompositions (Zhang and Bareinboim, 2018b; Plečko and Bareinboim, 2024) (this connection is explored explicitly in Appendix C). With the above example in mind, we list below the key contributions of this manuscript:
@@
    We develop an algorithm for attributing path-specific excess losses to different causal paths (Alg. 1), allowing the system designer to explain how the total excess loss is affected by different fairness constraints. In this context, we show the equivalence of Alg. 1 with a Shapley value (Shapley et al., 1953) approach (Prop 3),

### v2

@@ 1 Introduction
Automated decision-making systems based on machine learning and artificial intelligence are now commonly implemented in various critical sectors of society such as hiring, university admissions, law enforcement, credit assessments, and health care (Khandani, Kim, and Lo 2010; Mahoney and Mohen 2007; Brennan, Dieterich, and Ehret 2009). These technologies now significantly influence the lives of individuals and are frequently used in high-stakes settings (Topol 2019; Berk et al. 2021; Taddeo and Floridi 2018). As these systems replace or augment human decision-making processes, concerns about fairness and bias based on protected attributes such as race, gender, or religion have become a prominent consideration in the ML literature. The available data used to train automated systems may contain past and present societal biases as an imprint and therefore has the potential to perpetuate or even exacerbate discrimination against protected groups. This is highlighted by reports on biases in systems for sentencing (Angwin et al. 2016), facial recognition (Buolamwini and Gebru 2018), online ads (Sweeney 2013; Datta, Tschantz, and Datta 2015), and system authentication (Sanburn 2015), among many others. Despite the promise of AI to enhance human decision-making, the reality is that these technologies can also reflect or worsen societal inequalities. As alluded to before, the issue does not arise uniquely from the usage of automated systems; human-driven decision-making has long been analyzed in a similar fashion. Evidence of bias in human decision-making is abundant, including studies on the gender wage gap (Blau and Kahn 1992, 2017) and racial disparities in legal outcomes (Sweeney and Haney 1992; Pager 2003). Therefore, without proper care about fairness and transparency of the new generation of AI systems, it is unclear what its impact will be on the historically discriminated groups.
@@
Within the growing literature on fair machine learning, a plethora of fairness definitions have been proposed. Commonly considered statistical criteria, among others, include demographic parity (independence (Darlington 1971)), equalized odds (separation (Hardt, Price, and Srebro 2016)), and calibration (sufficiency (Chouldechova 2017)). These definitions, however, have been shown as mutually incompatible (Barocas and Selbst 2016; Kleinberg, Mullainathan, and Raghavan 2016). Despite a number of proposals, there is still a lack of consensus on what the appropriate measures of fairness are, and how statistical notions of fairness could incorporate moral values of the society at large. For this reason, a number of works explored the causal approaches to fair machine learning (Kusner et al. 2017; Kilbertus et al. 2017; Nabi and Shpitser 2018; Zhang and Bareinboim 2018b, a; Wu et al. 2019; Chiappa 2019; Plečko and Meinshausen 2020), and an in-depth discussion can be found in (Plečko and Bareinboim 2024). The main motivation for doing so is that the causal approach may allow the system designers to attribute the observed disparities between demographic groups to the causal mechanisms that underlie and generate them in the first place. In this way, by isolating disparities transmitted along different causal pathways, one obtains a more fine-grained analysis, and the capability to decide which causal pathways are deemed as unfair or discriminatory. More fundamentally, such considerations also form the basis of the legal frameworks for assessing discrimination in the United States and Europe. For instance, in the context of employment law, the disparate impact doctrine within the Title VII of the Civil Rights Act of 1964 (Act 1964) disallows any form of discrimination that results in a too large of a disparity between groups of interest. A core aspect of this doctrine, however, is the notion of business necessity (BN) or job-relatedness. Considerations of business necessity may allow variables correlated with the protected attribute to act as a proxy, and the law does not necessarily prohibit their usage due to their relevance to the business itself (or more broadly the utility of the decision-maker). Often, the wording that is used is that to argue business necessity in front of a court of law, the plaintiff needs to demonstrate that “there is no practice that is less discriminatory and achieves the same utility” (Els 1993). This concept illustrates the tension between fairness and utility, and demonstrates that we cannot be oblivious to considerations of utility from a legal standpoint. Moreover, demonstrating that a sufficient loss in accuracy results from imposing a fairness constraint has been previously used to justify business necessity considerations in some rulings of the European Court of Justice (Adams-Prassl, Binns, and Kelly-Lyth 2023; Weerts et al. 2023), emphasizing the importance of the topic studied in this paper.
Related work.
We mention three parts of related literature. First, we mention the literature exploring fairness-utility trade-offs, such as (Corbett-Davies et al. 2017). The essential argument is that an unconstrained predictor always achieves a greater or equal utility than a constrained one. Many works find that introducing fairness constraints reduces utility (Mitchell et al. 2021), and propose ways of handling the fairness-utility trade-off (Fish, Kun, and Lelkes 2016).
@@
However, other works in this literature still seem to be divided on the issue of whether trade-offs exist. For instance, (Rodolfa, Lamba, and Ghani 2021) finds that fairness and utility trade-offs are negligible in practice, while others argue that such trade-offs need not even exist (Maity et al. 2020; Dutta et al. 2020). Naturally, the implications on the predictor’s utility will strongly depend on the exact type of the fairness constraint that is enforced, and works that do not find a trade-off often focus on equality of odds (Hardt, Price, and Srebro 2016) ($\widehat{Y}{\perp\!\!\!\perp}X\mid Y$) or (multi)calibration (Chouldechova 2017) ($Y{\perp\!\!\!\perp}X\mid\widehat{Y}$). Notably, the former metric always allows for the perfect predictor $\widehat{Y}=Y$, and thus in settings with good predictive power, the cost of enforcing this constraint may indeed be negligible. The latter metric allows for the $L_{2}$-optimal prediction score, and improving miscalibration may sometimes also yield improvements in utility by serving as a type of regularization.
@@
Finally, in the causal fairness literature, the tension between fairness and utility has been largely unexplored. Some exceptions include works such as (Nilforoshan et al. 2022) which shows that for a decision policy that satisfies a causal fairness constraints, it is almost always possible to find another decision that has a higher utility and the same total variation (TV) measure. (Plecko and Bareinboim 2024a) performs a causal explanation on a decision score used for constructing a policy, and discusses how disparities in the decision score may influence utility. Our main aim of this paper is to fill in this gap, and provide a systematic way of analyzing the fairness-accuracy trade-off from a causal lens, and show that fairness and utility are almost always in a trade-off.
1.1 Motivating Example
We illustrate our approach in a simple linear setting:
@@ 1.1 Motivating Example
Consider variables $X,W,Y$ behaving according to the following linear system of equations:
@@
where $\epsilon_{w}\sim N(0,\sigma_{w}^{2}),\epsilon_{y}\sim N(0,\sigma_{y}^{2})$. Variable $X$ is the protected attribute, and $Y$ is the outcome of interest. The causal diagram of Eqs. 1-3 is shown in Fig. 2 (with the $Z$ set empty). Attribute $X$ can influence $Y$ along two different pathways: the direct path $X\to Y$, and the indirect path $X\to W\to Y$. Therefore, for considering fairness, we focus on fair linear predictors $\widehat{Y}^{S}$ of the form
$\widehat{Y}^{S}=\hat{\alpha}_{S}X+\hat{\gamma}_{S}W,$ (4)
where the predictor $\widehat{Y}^{S}$ removes effects in the set $S$, with $S$ ranging in {,DE,IE,{DE, $,\{\text{DE, IE}\}\}$ , { DE, IE } } (DE, IE stand for direct and indirect effects, and any subset of these could be removed). For instance, the optimal predictor $\widehat{Y}^{\emptyset}$, which is not subject to any fairness constraints, has the coefficients
$\hat{\alpha}_{\emptyset}=\alpha,\hat{\gamma}_{\emptyset}=\gamma,$ (5)
which are the ordinary least squares (OLS) coefficients. Therefore, its mean-squared error (MSE) can be computed as $\mathds{E}[Y-\widehat{Y}^{\emptyset}]^{2}=\sigma_{y}^{2}$. The DE-fair predictor $\widehat{Y}^{\text{DE}}$, which has the direct effect constrained to zero, has coefficients
$\hat{\alpha}_{\text{DE}}=0,\hat{\gamma}_{\text{DE}}=\gamma.$ (6)
The fully-fair predictor, labeled $\widehat{Y}^{\{\text{DE, IE}\}}$, has both direct and indirect effects constrained to zero, and thus has coefficients
$\hat{\alpha}_{\{\text{DE, IE}\}}=0,\hat{\gamma}_{\{\text{DE, IE}\}}=0.$ (7)
Thus, the corresponding MSE values for $\widehat{Y}^{\text{DE}}$ and $\widehat{Y}^{\{\text{DE, IE}\}}$ can be computed as (see Appendix A for computation details):
$\mathds{E}[Y-\widehat{Y}^{\text{DE}}]^{2}$ $=\sigma_{y}^{2}+\frac{\alpha^{2}}{2},$ (8)
@@
$\mathds{E}[Y-\widehat{Y}^{\{\text{DE, IE}\}}]^{2}$ $=\sigma_{y}^{2}+\frac{\alpha^{2}+\gamma^{2}\beta^{2}+\alpha\gamma\beta}{2}+\gamma^{2}\sigma_{w}^{2}.$ (9)
@@
Our goal is to decompose the total excess loss (TEL) originating from imposing the fairness constraints, defined as:
@@
$\text{TEL}:=\underbrace{\mathds{E}[Y-\widehat{Y}^{\{\text{DE, IE}\}}]^{2}}_{\text{fully-fair predictor's loss}}-\underbrace{\mathds{E}[Y-\widehat{Y}^{\emptyset}]^{2}}_{\text{unconstrained loss}}.$ (10)
TEL measures the excess loss (in terms of the increase in the MSE, compared to the unconstrained predictor) originating from the removal of the direct and indirect effects. This quantifies the excess loss traded off for enforcing fairness constraints. Our goal is to decompose the TEL to obtain path-specific contributions originating from the removal of direct and indirect effects as follows:
@@
TEL $=\underbrace{\mathds{E}[Y-\widehat{Y}^{\text{DE}}]^{2}-\mathds{E}[Y-\widehat{Y}^{\emptyset}]^{2}}_{\text{Term I = excess DE loss}}$ (11)
$+\underbrace{\mathds{E}[Y-\widehat{Y}^{\{\text{DE, IE}\}}]^{2}-\mathds{E}[Y-\widehat{Y}^{\text{DE}}]^{2}}_{\text{Term II = excess IE loss}}$ (12)
@@
$=\underbrace{\frac{\alpha^{2}}{2}}_{\text{Term I}}+\underbrace{\frac{\gamma^{2}\beta^{2}+\alpha\gamma\beta}{2}+\gamma^{2}\sigma_{w}^{2}}_{\text{Term II}}.$ (13)
Term I is the direct effect excess loss, incurred by constraining the direct effect to 0. Term II is the indirect effect excess loss, incurred by constraining the indirect effect to 0.
At the same time, enforcing fairness constraints may reduce the disparity between groups. For any predictor $\widehat{Y}$, we can measure the disparity using the difference in conditional expectations, called the total variation measure, defined as $\text{TV}_{x_{0},x_{1}}(\widehat{y})=\mathds{E}[\widehat{Y}\mid x_{1}]-\mathds{E}[\widehat{Y}\mid x_{0}]$. This measure is sometimes also referred to as the parity gap. Similarly as for the TEL, we compute the decrease in group disparity associated with removing DE and IE, by comparing the TV measures of the fully-fair $\widehat{Y}^{\{\text{DE, IE}\}}$ and unconstrained $\widehat{Y}^{\emptyset}$, and computing the TV difference (TVD, for short), defined as:
TVD $=\text{TV}(\widehat{Y}^{\{\text{DE, IE}\}})-\text{TV}(\widehat{Y}^{\emptyset})$ (14)
$=\underbrace{(\mathds{E}[\widehat{Y}^{\{\text{DE, IE}\}}\mid x_{1}]-\mathds{E}[\widehat{Y}^{\{\text{DE, IE}\}}\mid x_{0}])}_{\text{disparity after removing DE, IE}}$ (15)
$\quad-\underbrace{(\mathds{E}[\widehat{Y}^{\emptyset}\mid x_{1}]-\mathds{E}[\widehat{Y}^{\emptyset}\mid x_{0}])}_{\text{disparity before removing DE, IE}}=-\alpha-\beta\gamma.$ (16)
The TVD metric again decomposes into the contributions along direct and indirect effects:
TVD $=\underbrace{\text{TV}(\widehat{Y}^{\text{DE}})-\text{TV}(\widehat{Y}^{\emptyset})}_{\text{Term A = disparity reduction of DE}}$ (17)
$\quad+\underbrace{\text{TV}(\widehat{Y}^{\{\text{DE, IE}\}})-\text{TV}(\widehat{Y}^{\text{DE}})}_{\text{Term B = disparity reduction of IE}}.$ (18)
Terms A and B can be computed as:
[…]
$-(\mathds{E}[\widehat{Y}^{\emptyset}\mid x_{0}]-\mathds{E}[\widehat{Y}^{\emptyset}\mid x_{0}])=-\alpha$ (20)
[…]
$-(\mathds{E}[\widehat{Y}^{\text{DE}}\mid x_{0}]-\mathds{E}[\widehat{Y}^{\text{DE}}\mid x_{0}])=-\beta\gamma.$ (22)
Based on the excess loss and the reduction in disparity resulting from constraining a causal path to zero, we can quantify the fairness/utility trade-off through a causal lens. Prototypical instances for predictors $\widehat{Y}^{S}$ are visualized in Fig. 1, for three randomly drawn triples $(\alpha,\beta,\gamma)$. Predictor $\widehat{Y}^{\emptyset}$ is optimal and thus always has 00 excess loss (and hence lies on the vertical axis). Fully-fair predictor $\widehat{Y}^{\{\text{DE, IE}\}}$ removes both direct and indirect effects, and thus always has the TV measure equal to 00. Therefore, $\widehat{Y}^{\{\text{DE, IE}\}}$ always lies on the horizontal axis (corresponding to TV = 0). The slopes in the plot between $\widehat{Y}^{\emptyset}$, $\widehat{Y}^{\text{DE}}$, and $\widehat{Y}^{\text{DE, IE}}$ (indicated by arrows), geometrically capture the tension between excess loss and reducing discrimination upon imposing a constraint. These slopes are computed as the ratio
TV differenceExcess $\frac{\text{TV difference}}{\text{Excess Loss}},$ (23)
a quantity that we call the causal fairness-utility ratio (CFUR). Based on Eqs. 11-12 and 19-22 we can compute
CFUR $(\text{DE})=-\frac{1}{\alpha},$ (24)
$\text{CFUR}(\text{IE})=$ $-\frac{2\beta\gamma}{\gamma^{2}\beta^{2}+\alpha\gamma\beta+2\gamma^{2}\sigma_{w}^{2}},$ (25)
summarizing the fairness-utility trade-off for each path.
Figure 1: Total Variation (TV) measures vs. Excess Loss. Different colors represent trajectories obtained for different randomly sampled linear SCMs.
@@
The above example illustrates how in the simple linear case we can attribute the increased loss from imposing fairness constraints to the specific causal pathway in question. It also shows we can compute the associated change in the disparity between groups, quantified by the TV measure $\mathds{E}[\widehat{Y}\mid x_{1}]-\mathds{E}[\widehat{Y}\mid x_{0}]$. In this paper, we generalize the approach from the above example to a non-parametric setting, with the following key contributions:
@@
    We develop an algorithm for attributing path-specific excess losses to different causal paths (Alg. 1), allowing the system designer to explain how the total excess loss is affected by different fairness constraints. In this context, we show the equivalence of Alg. 1 with a Shapley value (Shapley et al. 1953) approach (Prop 3),